OpenStudy (anonymous):

If an object is dropped from a height of 144 ft, the function h(t) = –16t² + 144 gives the height of the object after t seconds. When will the object hit the ground?

OpenStudy (anonymous):

1.5 s 3 s 6 s 9 s

OpenStudy (bahrom7893):

–16t² + 144=0 -16t^2=-144 t=3